I fenced off a space in my back yard for a new puppy .the perimeter of the rectangular space is 6 meters if two opposite sides are each 2meters long what are the lengths of each of the other two sids
Answer by thepianist25(8)

A rectangle uses the perimeter formula P = 2l + 2w, where l is the length of one side, and w is the width of one side. Since this rectangle has a perimeter of 6 meters, and the two opposite sides (in this case, "l")are each 2 meters long, then our goal is to solve for w, the width:
P = 2l + 2w
(6) = 2(2) + 2w
6 = 4 + 2w
6 - 4 = 4 - 4 + 2w
2 = 2w
2/2 = 2w/2
1 = w, or w = 1
So, the width of each side is 1 meter.